OpenStudy (anonymous):

Let f(x) = 2x2 – 2x + 7 and g(x) = x2 + 9x – 6. Find f(x) – g(x). x2 – 11x + 13 x2 – 11x + 1 x2 + 7x + 1 x2 + 7x + 13

OpenStudy (rulnick):

x2-11x+13

OpenStudy (rulnick):

... 'cause 2x2-x2=x2, -2x-9x=-11x, and 7-(-6)=13

OpenStudy (anonymous):